# Asymmetric Graph Error Control with Low Complexity in Causal Bandits

### v1

@@ Appendix C Proof of Lemma 1
where the covariance matrix can be decomposed as $\bm{\Phi}_{\bm{a}}^{t}(i)=\bm{Q}_{i}\bm{\Lambda}_{i}\bm{Q}_{i}^{T}$. With the orthonormal matrix $\bm{Q}_{i}$, we have
$\left\lVert\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\right\rVert_{2}=\left\lVert\bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\right\rVert_{2},\ \bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\sim\mathcal{N}(\bm{0},\bm{\Lambda}_{i}).$ (79)
Since the squared norm of $\bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}$ is a weighted sum of independent squared Gaussian variables, it can be bounded from above as
$\left\lVert\bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\right\rVert_{2}^{2}\leq\lambda_{\max}(\bm{\Lambda}_{i})\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}\bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}$ (80)
$\hskip 60.0pt=\lambda_{\max}(\bm{\Phi}_{\bm{a}}^{t}(i))\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}\bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}.$ (81)
Notice that $\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}\bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}$ follows a Chi-squared distribution with $N$ degrees of freedom, with the $m$-th moment defined as
$\mathbb{E}\left[\left(\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}\bm{Q}_{i}^{T}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}\right)^{m}\right]=2^{m}\frac{\Gamma(m+N/2)}{\Gamma(N/2)},$ (82)
@@ Appendix D Proof of Lemma 2
$\bm{H}(\Delta\bm{B}_{\bm{a}}^{t})\doteq\begin{bmatrix}\bm{0}&\Delta\bm{B}_{\bm{a}}^{t}\\ \left(\Delta\bm{B}_{\bm{a}}^{t}\right)^{T}&\bm{0}\end{bmatrix}.$ (90)
@@
$\bm{H}(\Delta\bm{B}_{\bm{a}}^{t})=\sum_{i=1}^{N}\bm{H}_{i},\quad\bm{H}_{i}\doteq\begin{bmatrix}\bm{0}&\left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}\\ \left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}^{T}&\bm{0}\end{bmatrix},$ (92)
@@
$\bm{H}_{i}^{2}=\begin{bmatrix}\left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}\left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}^{T}&\bm{0}\\ \bm{0}&(\Delta\bm{B}_{\bm{a}}^{t})_{i}^{T}\left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}\end{bmatrix},$ (93)
@@ Appendix E Proof of Theorem 1
$\Delta\bm{\mu}_{\bm{a}}^{t}$ $=-\bm{\mu}_{\bm{a}}+(\bm{I}-\bm{B}_{\bm{a}})^{-T}\bm{\nu}+$
$\hskip 45.0pt(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-T}(\Delta\bm{B_{a}}^{t})^{T}(\bm{I}-\bm{B_{a}})^{-T}\bm{\nu}$ (107)
$=(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-T}(\Delta\bm{B_{a}}^{t})^{T}\bm{\mu}_{\bm{a}},$ (108)
@@
$\left|\left[\Delta\bm{\mu}_{\bm{a}}^{t}\right]_{N}\right|=\left|\bigl{[}(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}^{T}(\Delta\bm{B_{a}}^{t})^{T}\bm{\mu}_{\bm{a}}\right|$ (109)
$\hskip 24.0pt\leq\left\lVert\bigl{[}(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}\right\rVert_{2}\cdot\sigma_{\max}(\Delta\bm{B_{a}}^{t})\cdot\left\lVert\bm{\mu}_{\bm{a}}\right\rVert_{2}.$ (110)

### v2

@@ Appendix C Proof of Lemma 1
where the covariance matrix can be decomposed as $\bm{\Phi}_{\bm{a}}^{t}(i)=\bm{Q}_{i}\bm{\Lambda}_{i}{\bm{Q}}^{\top}_{i}$. With the orthonormal matrix $\bm{Q}_{i}$, we have
$\left\lVert\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\right\rVert_{2}=\left\lVert{\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\right\rVert_{2},\ {\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\sim\mathcal{N}(\bm{0},\bm{\Lambda}_{i}).$ (81)
Since the squared norm of ${\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}$ is a weighted sum of independent squared Gaussian variables, it can be bounded from above as
$\left\lVert{\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\right\rVert_{2}^{2}\leq\lambda_{\max}(\bm{\Lambda}_{i})\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}{\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}$ (82)
$\hskip 60.0pt=\lambda_{\max}(\bm{\Phi}_{\bm{a}}^{t}(i))\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}{\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}.$ (83)
Notice that $\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}{\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}$ follows a Chi-squared distribution with $N$ degrees of freedom, with the $m$-th moment defined as
$\mathbb{E}\left[\left(\bigl{\lVert}\bm{\Lambda}_{i}^{-1/2}{\bm{Q}}^{\top}_{i}\left[\Delta\bm{B}_{\bm{a}}^{t}\right]_{i}\bigr{\rVert}_{2}^{2}\right)^{m}\right]=2^{m}\frac{\Gamma(m+N/2)}{\Gamma(N/2)},$ (84)
@@ Appendix D Proof of Lemma 2
$\bm{H}(\Delta\bm{B}_{\bm{a}}^{t})\doteq\begin{bmatrix}\bm{0}&\Delta\bm{B}_{\bm{a}}^{t}\\ {\left(\Delta\bm{B}_{\bm{a}}^{t}\right)}^{\top}&\bm{0}\end{bmatrix}.$ (92)
@@
$\bm{H}(\Delta\bm{B}_{\bm{a}}^{t})=\sum_{i=1}^{N}\bm{H}_{i},\quad\bm{H}_{i}\doteq\begin{bmatrix}\bm{0}&\left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}\\ {\left(\Delta\bm{B}_{\bm{a}}^{t}\right)}^{\top}_{i}&\bm{0}\end{bmatrix},$ (94)
@@
$\bm{H}_{i}^{2}=\begin{bmatrix}\left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}{\left(\Delta\bm{B}_{\bm{a}}^{t}\right)}^{\top}_{i}&\bm{0}\\ \bm{0}&{(\Delta\bm{B}_{\bm{a}}^{t})}^{\top}_{i}\left(\Delta\bm{B}_{\bm{a}}^{t}\right)_{i}\end{bmatrix},$ (95)
@@ Appendix E Proof of Theorem 1
$\Delta\bm{\mu}_{\bm{a}}^{t}$ $=-\bm{\mu}_{\bm{a}}+{(\bm{I}-\bm{B}_{\bm{a}})}^{-\top}\bm{\nu}+$
$\hskip 45.0pt{(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})}^{-\top}{(\Delta\bm{B_{a}}^{t})}^{\top}{(\bm{I}-\bm{B_{a}})}^{-\top}\bm{\nu}$ (109)
$={(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})}^{-\top}{(\Delta\bm{B_{a}}^{t})}^{\top}\bm{\mu}_{\bm{a}},$ (110)
@@
$\left|\left[\Delta\bm{\mu}_{\bm{a}}^{t}\right]_{N}\right|=\left|{\bigl{[}(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}}^{\top}_{N}{(\Delta\bm{B_{a}}^{t})}^{\top}\bm{\mu}_{\bm{a}}\right|$ (111)
$\hskip 24.0pt\leq\left\lVert\bigl{[}(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}\right\rVert_{2}\cdot\sigma_{\max}(\Delta\bm{B_{a}}^{t})\cdot\left\lVert\bm{\mu}_{\bm{a}}\right\rVert_{2}.$ (112)